OpenStudy (anonymous):

Find the equation of the line passing through (0,−1) and perpendicular to the line passing through the points (1, 2) and (5, 4)

OpenStudy (anonymous):

2x+y+1=0

OpenStudy (anonymous):

what's this?

OpenStudy (anonymous):

the equation u want

OpenStudy (anonymous):

slope of the line through (1,2) and (5,4) is \[\frac{4-2}{5-1}=\frac{2}{4}=\frac{1}{2}\]

OpenStudy (anonymous):

how did u get the equation?

OpenStudy (anonymous):

slope of perpendicular line is therefore -2 you already have the y - intercept so equation is \[y=-2x-1\]

OpenStudy (anonymous):

ok. at first solve the equation of the line passing through (1,2) (5,4).do u know how to do that?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

now.lthe multiplication of slopes of 2 lines perpendicular to each other is=-1

OpenStudy (anonymous):

so,u get the slope of the line.u know that y=mx+c.put the values of y,m & x.u get c.so,ur equation is ready